# Transition of Anticoncentration in Gaussian Boson Sampling

### 13 Dec 2023 v1

@@ Acknowledgements.
S4 Details on Definitions of Anticoncentration
@@
In this section, we discuss in somewhat more detail the relevance of anticoncentration to the argument for hardness of sampling from the output distribution of Gaussian Boson Sampling (GBS). This argument makes use of an approximate counting algorithm due to Stockmeyer (1983). Roughly, we assume that there is an efficient sampling algorithm for GBS that, given a linear-optical unitary $U$, samples from a distribution $Q_{U}$ which is close up to a constant $\epsilon>0$ to the ideal GBS distribution $P_{U}$ (recall Eq. 1 of the main text) in total-variation distance
@@
The relevant quantity of interest to anticoncentration is thus the inverse average collision probability $p_{2}(\mathrm{U}(m))=1/P_{2}(\mathrm{U}(m)$. Because by hiding the first moment $\operatorname*{\mathbb{E}}_{U}[P_{U}(\mathbf{n})]$ must evaluate to the inverse size of the effective sample space, we can rewrite $p_{2}$ for GBS as
@@
    And we say that it does not anticoncentrate if $p_{2}=O(1/\exp(n))$.
@@
Ignoring the average over unitaries, $p_{2}$ upper-bounds the support of the distribution by $p_{2}|\Omega|$, as the maximum-entropy state is the uniform distribution. Hence, an exponentially small value of $p_{2}$ implies that the average support of the outcome distributions $P_{U}$ is exponentially small, implying that at least a constant fraction (over $U$) of the distributions $P_{U}$ has exponentially small support, and conversely exponentially larger than uniform probabilities on that support. At least for those distributions, this implies an exponentially larger error tolerance compared to $1/|\Omega|$. Such an exponentially larger error tolerance makes the approximate average-case hardness conjecture significantly stronger, presumably even untenable.
@@
We argue that this is a meaningful regime in the sense that there is a stronger—but not inconceivable—approximate average-case hardness conjecture associated with the weak anticoncentration regime. To see this, observe that weak anticoncentration implies anticoncentration in the sense of Eq. S38 with $\gamma(\alpha)=\Omega(1/\textrm{poly}(n))$, which means that an inverse polynomial fraction of the outcome probabilities are larger than uniform. Technically, using Stockmeyer’s algorithm we can thus achieve a multiplicative error for an inverse polynomial fraction of the outcome probabilities. To rule out an efficient classical sampler, we thus need to conjecture approximate average-case hardness with constant relative errors for any inverse polynomial fraction of the instances. Equivalently, we can formulate a similar conjecture for a polynomially large relative or subexponentially large additive error on a constant fraction. While clearly much stronger than the requirement of anticoncentration, this is qualitatively different from the lack of anticoncentration scenario (NA), where the difference is exponential.
S5 Scattershot Boson Sampling Explanation of the Transition in Anticoncentration
@@
Let us for simplicity assume that the photon numbers observed in the postselection measurement are Poissonian distributed as $p(l)=\sinh^{2l}(r)\mathrm{e}^{-\sinh^{2}(r)}/l!$ with mean $\sinh^{2}(r)=n/k$. Then the probability of observing a collision scales as
$\Pr[\text{collision}]=1-\left(\left(1+\frac{n}{k}\right)e^{-n/k}\right)^{k}=1-\exp\left(-\frac{n^{2}}{2k}+\sum_{l=3}^{\infty}\frac{(-1)^{l+1}}{l}\frac{n^{l}}{k^{l-1}}\right),$ (S44)
assuming $k=\omega(n)$. This collision probability remains lower bounded by a constant for $k=O(n^{2})$, but vanishes for any $k=\omega(n^{2})$. Thus, the probability of a collision in the input state of SBS remains high until $k=\Theta(n^{2})$ and decays then. But because in SBS the roles of the (postselected) input state and the output state are symmetric, a collision implies a failure of hiding and therefore a failure of anticoncentration in the regime $k=O(n^{2})$. Conversely, for $k=\omega(n^{2})$ we believe that hiding holds Jiang (2006); Aaronson and Arkhipov (2013), and hence Lemma 8.8 of Aaronson and Arkhipov (2013) shows weak anticoncentration for SBS with the inverse average collision probability $p_{2}=1/n$ in this regime.

### 20 Mar 2024 v2

@@ Acknowledgements.
S4 Approximate Hiding and Asymptotics of the First Moment
In this section, we discuss more thoroughly the connection between hiding, the relevant sample space, and the first moment of squared hafnians of generalized COE matrices.
In the main text, we introduce the normalized average outcome-collision probability as a measure of anticoncentration. Fixing the output state to have $2n$ photons, we write this as $|\Omega_{2n}|\operatorname*{\mathbb{E}}_{U\in U(m)}[\sum_{\mathbf{n}\in\Omega_{2n}}P_{U}(\mathbf{n})^{2}]$, where $\Omega_{2n}$ is the space of collision-free outcomes with $2n$ photons in $m$ modes, and its size, which we write as $|\Omega_{2n}|$, is simply $\binom{m}{2n}$. We here work specifically with the non-collisional sample space because, in order for hiding to hold, collisions have to be negligible (a non-negligible likelihood of repeated columns in $U^{\top}_{1_{k},\mathbf{n}}U_{1_{k},\mathbf{n}}$ would prevent this distribution from being well approximated by $X^{\top}X$ with $X$ Gaussian). And, indeed, when $n=o(\sqrt{m})$, it is easy to see that the size of the full sample space of $2n$ photons in $m$ modes, $\binom{m+2n-1}{2n}$, approaches $|\Omega_{2n}|=\binom{m}{2n}$ when $n\gg 1$. In particular,
$\frac{m^{2n}}{(2n)!}\leq\frac{(m+2n-1)!}{(m-1)!(2n)!}\leq\frac{(m+2n-1)^{2n}}{(2n)!}$ (S36)
and
$\frac{\frac{(m+2n-1)^{2n}}{(2n)!}}{\frac{m^{2n}}{(2n)!}}=\left(1+\frac{2n-1}{m}\right)^{2n}\stackrel{{\scriptstyle n\gg 1}}{{\longrightarrow}}1$ (S37)
because $(2n-1)/m=o(1/n)$. That is, $\Omega_{2n}$ is the dominant contribution to the full sample space.
We proceed to then replace $|\Omega_{2n}|$ with the expected value of the outcome probabilties, $\mathbb{E}_{U}[P_{U}(\mathbf{n})]$, that is, the first moment over input unitaries of a specific outcome. This holds assuming that the hiding property in Conjecture 1 holds. Roughly, hiding ensures that we do not preference any individual outcome, meaning we can replace the expected value over all probabilities with that over unitaries for a single probability. By linearity of expectation (and the fact that probabilities sum to unity), this expectation over unitaries should simply be the inverse of the size of the sample space of non-collisional outcomes. Finally, Conjecture 1 also gives us an approximate equality between $\mathbb{E}_{U}[P_{U}(\mathbf{n})]$ and $M_{1}$, the first moment of the squared hafnian of generalized COE matrices (properly rescaled to contain the correct prefactors).
We therefore now show that our calculation of the first moment in the hiding regime is consistent with the above discussion in the sense that $\mathbb{E}_{U}[P_{U}(\mathbf{n}|\sum_{i}\mathbf{n}_{i}=2n)]=\mathbb{E}_{U}[P_{U}(\mathbf{n})]/P(2n)$ is asymptotically equal to $|\Omega_{2n}|^{-1}=\binom{m}{2n}^{-1}$ assuming Conjecture 1. Here, $P(2n)$ is the probability that our output is in the $2n$-photon sector (i.e., the probability that $\Omega_{2n}$ is the proper sample space to consider in the first place).
Recall our input state has the first $k$ of $m$ modes prepared in the single-mode squeezed vacuum state with identical squeezing parameter $r$, and the remaining $m-k$ modes are prepared in the vacuum state. The probability of an outcome $\mathbf{n}$ is given by Eq. 1:
$P_{U}(\mathbf{n})=\frac{\tanh^{2n}r}{\cosh^{k}r}\absolutevalue{\mathrm{Haf}(U^{\top}_{1_{k},\mathbf{n}}U_{1_{k},\mathbf{n}})}^{2},$ (S38)
where $U_{1_{k},\mathbf{n}}$ is the submatrix of $U$ given by the first $k$ rows and the columns dictated by where $\mathbf{n}$ is nonzero. Define $\tilde{U}_{1_{k},\mathbf{n}}\coloneqq mU_{1_{k},\mathbf{n}}$. Using multiplicativity of the Hafnian, one finds
$P_{U}(\mathbf{n})=\frac{\tanh^{2n}r}{\cosh^{k}r}\frac{1}{m^{2n}}\absolutevalue{\mathrm{Haf}\left(\tilde{U}_{1_{k},\mathbf{n}}\tilde{U}^{\top}_{1_{k},\mathbf{n}}\right)}^{2}.$ (S39)
Assuming Conjecture 1, then $U^{\top}_{1_{k},\mathbf{n}}U_{1_{k},\mathbf{n}}\sim X^{\top}X$ where $X\sim\mathcal{N}(0,1/m)_{c}^{k\times 2n}$, which means $\tilde{U}^{\top}_{1_{k},\mathbf{n}}\tilde{U}_{1_{k},\mathbf{n}}\sim X^{\top}X$, but now $X\sim\mathcal{N}(0,1)_{c}^{k\times 2n}$. Then, by Conjecture 1 and Theorem 2, we find
$\underset{U\in\mathrm{U}(m)}{\operatorname*{\mathbb{E}}}\left[\absolutevalue{\mathrm{Haf}\left(\tilde{U}^{\top}_{1_{k},\mathbf{n}}\tilde{U}_{1_{k},\mathbf{n}}\right)}^{2}\right]\approx\underset{X\in\mathcal{G}^{k\times 2n}}{\operatorname*{\mathbb{E}}}\left[\absolutevalue{\mathrm{Haf}\left(X^{\top}X\right)}^{2}\right]=\frac{(2n)!}{2^{n}n!}\frac{(k+2n-2)!!}{(k-2)!!},$ (S40)
where the first part of the equation is not an equality precisely because the hiding in Conjecture 1 is not exact. This implies that
$\mathbb{E}_{U}[P_{U}(\mathbf{n})]\approx\frac{\tanh^{2n}r}{\cosh^{k}r}\frac{1}{m^{2n}}\frac{(2n)!}{2^{n}n!}\frac{(k+2n-2)!!}{(k-2)!!}.$ (S41)
Now, a single-mode squeezed vacuum state with squeezing parameter $r$ and phase $\phi$ has Fock-state expansion given by
|$\ket{\textrm{SMSV}}=\frac{1}{\sqrt{\cosh r}}\sum_{\ell=0}^{\infty}(-e^{i\phi}\tanh r)^{\ell}\frac{\sqrt{(2\ell)!}}{2^{\ell}\ell!}\ket{2\ell}.$ (S42)
Therefore, the probability of measuring $2\ell$ photons is
|$\absolutevalue{\braket{2\ell}{\textrm{SMSV}}}^{2}=\frac{\tanh^{2\ell r}}{\cosh r}\frac{(2\ell)!}{(2^{\ell}\ell!)^{2}}.$ (S43)
Given $k$ independent single-mode squeezed vacuum states, the probability of finding $2n$ total photons is the $k$-fold convolution of the Fock-basis probability distribution of one single-mode squeezed vaccuum state:
$P(2n)=\sum_{2\ell_{1}+\dots+2\ell_{k}=2n}\prod_{i=1}^{k}\frac{\tanh^{2\ell_{i}r}}{\cosh r}\frac{(2\ell_{i})!}{(2^{\ell_{i}}\ell_{i}!)^{2}}=\frac{\tanh^{2n}r}{\cosh^{k}r}\frac{1}{2^{2n}}\sum_{2\ell_{1}+\dots+2\ell_{k}=2n}\prod_{i=1}^{k}\binom{2\ell_{1}}{\ell_{1}}.$ (S44)
This probability distribution is unchanged if the $k$ independent single-mode squeezed vacuum states are acted upon by a linear-optical unitary before measurement (such a unitary does not change the photon number, only the location of the photons). The combinatorial identity at the core of this $k$-fold convolution has been calculated before in Refs. Li and Wenchang Chu (2019); Chang and Xu (2011). Specifically,
$\sum_{2\ell_{1}+\dots+2\ell_{k}=2n}\prod_{i=1}^{k}\binom{2\ell_{1}}{\ell_{1}}=4^{n}\binom{n-1+k/2}{n},$ (S45)
where we note that Eq. S45 holds even in the case where $k$ is odd using a generalization of the binomial coefficients in terms of the $\Gamma$ function.
The overall probability of finding $2n$ photons from $k$ independent single-mode squeezed vacuum states, even after the application of a linear optical unitary, is therefore
$P(2n)=\frac{\tanh^{2n}r}{\cosh^{k}r}\frac{1}{2^{2n}}4^{n}\binom{n-1+k/2}{n}=\frac{\tanh^{2n}r}{\cosh^{k}r}\binom{n-1+k/2}{n}.$ (S46)
We note that this expression, but not the full derivation, is also provided in Ref. Kruse (2019). A bit of algebraic manipulation reveals
$P(2n)=\frac{\tanh^{2n}r}{\cosh^{k}r}\binom{n-1+k/2}{n}=\frac{\tanh^{2n}r}{\cosh^{k}r}\frac{(2n-1)!!(k+2n-2)!!}{(2n)!(k-2)!!}=\frac{1}{(2n)!}\frac{\tanh^{2n}r}{\cosh^{k}r}\underset{X\sim\mathcal{G}^{k\times 2n}}{\operatorname*{\mathbb{E}}}\left[\absolutevalue{\mathrm{Haf}(X^{\top}X)}^{2}\right].$ (S47)
According to Eq. S41, then
$P(2n)\approx\frac{m^{2n}}{(2n)!}\mathbb{E}_{U}[P_{U}(\mathbf{n})],$ (S48)
which finally implies
$\frac{\mathbb{E}_{U}[P_{U}(\mathbf{n})]}{P(2n)}\approx\frac{(2n)!}{m^{2n}}\approx\binom{m}{2n}^{-1}=|\Omega_{2n}|^{-1},$ (S49)
where the first approximation is due to the fact that hiding is not exact, and the second approximation holds in the photon non-collisional regime.
S5 Details on Definitions of Anticoncentration
In this section, we discuss some of the details behind our definition of anticoncentration and how it relates to the standard notion of anticoncentration often used in the literature. We also discuss how these different definitions interact when it comes to showing anticoncentration holds for the exact distribution of the output probabilities of GBS given anticoncentration of the approximate distribution.
S5.1 Definitions
@@
We first discuss in somewhat more detail the relevance of anticoncentration to the argument for hardness of sampling from the output distribution of Gaussian Boson Sampling (GBS). This argument makes use of an approximate counting algorithm due to Stockmeyer (1983). Roughly, we assume that there is an efficient sampling algorithm for GBS that, given a linear-optical unitary $U$, samples from a distribution $Q_{U}$ which is close up to a constant $\epsilon>0$ to the ideal GBS distribution $P_{U}$ (recall Eq. 1 of the main text) in total-variation distance
@@
The relevant quantity of interest to anticoncentration is thus the inverse average collision probability $p_{2}(\mathrm{U}(m))=1/P_{2}(\mathrm{U}(m)$. Because by hiding the first moment $\operatorname*{\mathbb{E}}_{U}[P_{U}(\mathbf{n})]$ must evaluate to the inverse size of the sample space, we can rewrite $p_{2}$ for GBS as
@@
    And we say that it does not anticoncentrate if $p_{2}=O(1/n^{a})$ for any constant $a>0$.
@@
Ignoring the average over unitaries, $p_{2}$ upper-bounds the support of the distribution by $p_{2}|\Omega|$, as the maximum-entropy state is the uniform distribution. Let us assume for simplicity that $p_{2}$ is actually exponentially small. An exponentially small value of $p_{2}$ implies that the average support of the outcome distributions $P_{U}$ is exponentially small, implying that at least a constant fraction (over $U$) of the distributions $P_{U}$ has exponentially small support, and conversely exponentially larger than uniform probabilities on that support. At least for those distributions, this implies an exponentially larger error tolerance compared to $1/|\Omega|$. Such an exponentially larger error tolerance makes the approximate average-case hardness conjecture significantly stronger, presumably even untenable.
@@
We argue that this is a meaningful regime in the sense that there is a stronger—but not inconceivable—approximate average-case hardness conjecture associated with the weak anticoncentration regime. To see this, observe that weak anticoncentration implies anticoncentration in the sense of Eq. S52 with $\gamma(\alpha)=\Omega(1/\textrm{poly}(n))$, which means that an inverse polynomial fraction of the outcome probabilities are larger than uniform. Technically, using Stockmeyer’s algorithm we can thus achieve a multiplicative error for an inverse polynomial fraction of the outcome probabilities. To rule out an efficient classical sampler, we thus need to conjecture approximate average-case hardness with constant relative errors for any inverse polynomial fraction of the instances. Equivalently, we can formulate a similar conjecture for a polynomially large relative or subexponentially large additive error on a constant fraction. While clearly much stronger than the requirement of anticoncentration, this is qualitatively different from the lack of anticoncentration scenario (NA), where the difference is superpolynomial.
S5.2 Anticoncentration of the Exact Distribution
We also need to show that our definition of anticoncentration allows us to translate between anticoncentration of the approximate distribution based on the hafnians of random Gaussian matrices, which we will refer to as $P_{X}(\mathbf{n})$, and anticoncentration of the true distribution, $P_{U}(\mathbf{n})$. For a given output $\mathbf{n}$, let $\mathcal{D}_{U}$ be the distribution of the symmetric product $U_{1_{k},\mathbf{n}}^{\top}U_{1_{k},\mathbf{n}}$ with $U\in\mathrm{U}(m)$. Let $\mathcal{D}_{X}$ be the distribution of the symmetric product $X^{\top}X$ with $X\sim\mathcal{N}(0,1/m)_{c}^{k\times 2n}$. In Conjecture 1, we conjecture that $\mathcal{D}_{U}$ and $\mathcal{D}_{X}$ become close in total variation distance when $n=o(\sqrt{m})$. However, precisely how close these two distributions are is crucial to whether or not anticoncentration translates between the two output probabily distributions. In what follows, we will refer to anticoncentration in the sense of Eq. S52 as “standard” anticoncentration, and our definition of anticoncentration as “moment-based.”
Ideally, we would be able to prove that statements about moment-based anticoncentration of $P_{X}(\mathbf{n})$ imply equivalent statements about moment-based anticoncentration of $P_{U}(\mathbf{n})$. However, under worst-case assumptions, we can only show that moment-based anticoncentration of $P_{X}(\mathbf{n})$ implies standard anticoncentration of $P_{U}(\mathbf{n})$. To understand this, let us fix some notation. Let also $\mathbbm{1}[\cdot]$ be an indicator function which is $1$ if the argument is true and 00 if it is false. Let $\mathrm{d}\mu$ be the Lebesgue measure on $\mathbb{C}^{2kn}$ (as we consider $k\times 2n$ complex matrices) and $p_{U}(A)$, $p_{X}(A)$ be the respective probabilities of generating $A$ from $\mathcal{D}_{U}$ and $\mathcal{D}_{X}$.
Now, let the total-variation distance between $\mathcal{D}_{U}$ and $\mathcal{D}_{X}$ be less than $\delta$. Then
$\Pr_{U\in\mathrm{U}(m)}[P_{U}(\mathbf{n})\geq\epsilon]$ $=\int\mathrm{d}\mu\,p_{U}(A)\mathbbm{1}[P_{A}(\mathbf{n})\geq\epsilon]$ (S57)
$=\int\mathrm{d}\mu\,(p_{U}(A)-p_{X}(A)+p_{X}(A))\mathbbm{1}[P_{A}(\mathbf{n})\geq\epsilon]$ (S58)
$=\int\mathrm{d}\mu\,(p_{U}(A)-p_{X}(A))\mathbbm{1}[P_{A}(\mathbf{n})\geq\epsilon]+\int\mathrm{d}\mu\,p_{X}(A)\mathbbm{1}[P_{A}(\mathbf{n})\geq\epsilon]$ (S59)
$\geq-2\delta+\Pr_{X\in\mathcal{G}}[P_{X}(\mathbf{n})\geq\epsilon].$ (S60)
In this calculation, we have used the Radon-Nikodym theorem Folland (2007) to express the probability measures that define $\mathcal{D}_{U}$ and $\mathcal{D}_{X}$ as $p_{U}(A)\mathrm{d}\mu$ and $p_{X}(A)\mathrm{d}\mu$, respectively. Therefore
$\Pr_{U\in\mathrm{U}(m)}\left[P_{U}(\mathbf{n})\geq\frac{\alpha}{|\Omega_{2n}|}\right]\geq\Pr_{X\in\mathcal{G}}\left[P_{X}(\mathbf{n})\geq\frac{\alpha}{|\Omega_{2n}|}\right]-2\delta\geq(1-\alpha)^{2}\frac{1}{m_{2}(k,n)}-2\delta.$ (S61)
The final step follows from the Paley-Zygmund inequality for the approximate distribution. This proves that we can translate statements on anticoncentration as long as $2\delta$ is smaller than $m_{2}(k,n)^{-1}$, which, as we show in the main text, means $\delta=o(n^{-1/2})$.
With this in mind, we can make the following more precise version of Conjecture 1 such that, if it holds, moment-based weak anticoncentration of the approximate distribution implies standard weak anticoncentration of the exact distribution:
Conjecture 1 (Formal).
Let $\mathcal{D}_{U}$ be the distribution of the symmetric product $U_{1_{k},\mathbf{n}}^{\top}U_{1_{k},\mathbf{n}}$ with $U$ unitary and $\mathbf{n}$ some non-collisional outcome of a Gaussian Boson Sampling experiment. Let $\mathcal{D}_{X}$ be the distribution of the symmetric product $X^{\top}X$ with $X\sim\mathcal{N}(0,1/m)_{c}^{k\times 2n}$. Then, for any $k$ such that $1\leq k\leq m$, and for any $\delta>0$ such that $m\geq n^{2}/\delta$,
$\mathsf{tvd}(\mathcal{D}_{U},\mathcal{D}_{X})=O(\delta).$ (S62)
Specifically, if $\delta=o(n^{-1/2})$, then $m\geq n^{5/2}$.
The motivation behind the choice of $m\geq n^{2}/\delta$ is based on the equivalent conjecture for Fock Boson Sampling in Ref. Aaronson and Arkhipov (2013). There, the authors are able to prove the equivalent result for $m\geq n^{5+\epsilon}/\delta$ (for arbitrarily small, constant $\epsilon$), but they suspect that the result can be pushed further to $m\geq n^{2}/\delta$. We note that this choice makes our formal conjecture slightly stronger than the equivalent formal conjecture in Ref. Deshpande et al. (2022a).
As we have shown, in order to translate our results on moment-based weak anticoncentration from the approximate to the true distribution in the worst case, we require $\delta=o(n^{-1/2})$. Therefore, in order to translate statements about anticoncentration, the formal version of our conjecture requires $m\geq n^{5/2}$.
However, it is worth noting that we do not believe that this worst-case scenario truly reflects the way in which $P_{X}(\mathbf{n})$ approaches $P_{U}(\mathbf{n})$, i.e., where all of the error is concentrated on a single probability. In general, the intuition is that if hiding holds, then it is more likely that the errors are more evenly distributed amongst all of the exponentially many output probabilities. Using this intuition, each individual probability only receives an error of approximately $\delta/|\Omega_{2n}|$. If this is true, then we can show that moment-based weak anticoncentration of $P_{X}(\mathbf{n})$ does actually imply the same for $P_{U}(\mathbf{n})$. Specifically, say that $P_{U}(\mathbf{n})\approx P_{X}(\mathbf{n})\pm\delta/|\Omega_{2n}|\approx P_{X}(\mathbf{n})\pm\delta\operatorname*{\mathbb{E}}[P_{X}(\mathbf{n})]$ (as per Section S4). Then
$\frac{\operatorname*{\mathbb{E}}[P_{U}(\mathbf{n})^{2}]}{(\operatorname*{\mathbb{E}}[P_{U}(\mathbf{n})])^{2}}$ $\approx\frac{\operatorname*{\mathbb{E}}[(P_{X}(\mathbf{n})\pm\delta\operatorname*{\mathbb{E}}[P_{X}(\mathbf{n})])^{2}]}{(\operatorname*{\mathbb{E}}[P_{X}(\mathbf{n})\pm\delta\operatorname*{\mathbb{E}}[P_{X}(\mathbf{n})]])^{2}}$ (S63)
$=\frac{\operatorname*{\mathbb{E}}[P_{X}(\mathbf{n})^{2}]\pm 2\delta\operatorname*{\mathbb{E}}[P_{X}(\mathbf{n})]^{2}+\delta^{2}\operatorname*{\mathbb{E}}[P_{X}(\mathbf{n})]^{2}}{(1\pm\delta)^{2}\operatorname*{\mathbb{E}}[P_{X}(\mathbf{n})]^{2}}$ (S64)
$\approx\frac{1}{(1\pm\delta)^{2}}\frac{\operatorname*{\mathbb{E}}[P_{X}(\mathbf{n})^{2}]}{(\operatorname*{\mathbb{E}}[P_{X}(\mathbf{n})])^{2}}+\frac{\pm 2\delta+\delta^{2}}{(1\pm\delta)^{2}}$ (S65)
$=\frac{1}{(1\pm\delta)^{2}}\frac{\operatorname*{\mathbb{E}}[P_{X}(\mathbf{n})^{2}]}{(\operatorname*{\mathbb{E}}[P_{X}(\mathbf{n})])^{2}}+1-\frac{1}{(1\pm\delta)^{2}}$ (S66)
$\leq\frac{1}{(1-\delta)^{2}}\frac{\operatorname*{\mathbb{E}}[P_{X}(\mathbf{n})^{2}]}{(\operatorname*{\mathbb{E}}[P_{X}(\mathbf{n})])^{2}}+1.$ (S67)
In our case, where the normalized second moment of $P_{X}(\mathbf{n})$ scales at least polynomially in $n$, and $\delta$ scales inverse polynomially in $n$, weak anticoncentration or lack of anticoncentration of $P_{X}(\mathbf{n})$ in terms of the normalized second moment adequately translates to $P_{U}(\mathbf{n})$ as well. Note that, in this case, we are assuming that $\delta$, which is the total variation distance between the distributions of matrices, extends to a bound on the total variation distance between the probabilities themselves. This intuitively arises from the fact that any map from the distribution of the matrices to probabilities must be bounded, meaning we can translate the total variation distance from one to the other (however, formalizing this would require dealing with some subtleties induced by the fact that the hafnian of a product of Gaussians is not technically bounded, but any large hafnians only arise with extremely small probabilities).
S6 Scattershot Boson Sampling Explanation of the Transition in Anticoncentration
@@
Recall that a two-mode squeezed vacuum state with squeezing parameter $r$ and phase $\phi$ has a Fock expansion given by
|$\ket{\mathrm{TMSV}}=\frac{1}{\cosh(r)}\sum_{\ell=0}^{\infty}(-e^{i\phi}\tanh r)^{n}\ket{nn},$ (S69)
thus leading to a probability of measuring $\ell$ photons in one mode of $\tanh^{2\ell}r/\cosh^{2}r$. Therefore, if the input consists of $k$ two-mode squeezed vacuum states, then the probability that, after measuring one half of each state, one observes a collision is
$\Pr[\text{collision}]=1-\left(\frac{1}{\cosh^{2}r}+\frac{\tanh^{2}r}{\cosh^{2}r}\right)^{k}=1-\left(\frac{1}{1+n/k}+\frac{n/k}{(1+n/k)^{2}}\right)^{k}=1-\left(1-\left(\frac{n/k}{1+n/k}\right)^{2}\right)^{k}.$ (S70)
We can rewrite this via Taylor series as
$\Pr[\text{collision}]=1-\exp\left(-\frac{n^{2}}{k}+kO(n/k)^{3})\right)$ (S71)
assuming $k=\omega(n)$. This collision probability remains lower bounded by a constant for $k=O(n^{2})$, but vanishes for any $k=\omega(n^{2})$. Thus, the probability of a collision in the input state of SBS remains high until $k=\Theta(n^{2})$ and decays then. But because in SBS the roles of the (postselected) input state and the output state are symmetric, a collision implies a failure of hiding and, therefore, a failure of anticoncentration in the regime $k=O(n^{2})$. Conversely, for $k=\omega(n^{2})$ we believe that hiding holds Jiang (2006); Aaronson and Arkhipov (2013), and hence Lemma 8.8 of Aaronson and Arkhipov (2013) shows weak anticoncentration for SBS with the inverse average collision probability $p_{2}=1/n$ in this regime.